The sum of the first n odd positive integers is 9409. What is the value of n ?
Answer by drk(1908)

Sum of 1 = 1 = 1*1 = 1^2
sum of 1 and 3 = 4 = 2*2 = 2^2
Sum of 1,3,5 = 9 = 3*3 = 3^2
and so on.
Sum of 1,3,5, . . . , n = n^2
So,
n^2 = 9409
n = 97